# Probing light sterile neutrinos in left-right symmetric models with displaced vertices and neutrinoless double beta decay

### v1

@@ 1 Introduction
Recently, there has been much interest in sterile neutrinos with GeV-scale masses (see, for instance, Refs. Atre:2009rg ; Bondarenko:2018ptm ; Helo:2013esa ; Drewes:2019fou ; Cottin:2022nwp ; Liu:2023gpt ), especially in the context of searches for long-lived particles (LLPs) (see Refs. Alimena:2019zri ; Lee:2018pag ; Curtin:2018mvb ; Beacham:2019nyx for recent reviews of LLPs). In the minimal scenario where the only interactions between the sterile neutrinos and the SM particles are mediated by the active-sterile neutrino mixings $U^{2}$, the current bounds on the sterile neutrinos in this mass range hint that these sterile neutrinos, if existent, would be long-lived Deppisch:2015qwa ; Bryman:2019bjg ; Abdullahi:2022jlv . Beam-dump, neutrino, and collider experiments produce mesons in very large numbers. These mesons can produce sterile neutrinos with masses in the MeV-GeV range through rare decays. With a long lifetime, these sterile neutrinos travel a macroscopic distance before potentially decaying inside a detector volume placed a few to a few hundred meters away. Past studies have investigated these minimal scenarios extensively and indicated that (proposed) future experiments such as DUNE Krasnov:2019kdc ; Ballett:2019bgd ; Berryman:2019dme ; Coloma:2020lgy ; Moghaddam:2022tac ; Plows:2022gxc ; Gunther:2023vmz , ANUBIS Hirsch:2020klk ; DeVries:2020jbs , SHiP SHiP:2015vad ; Alekhin:2015byh ; SHiP:2018xqw ; DeVries:2020jbs , and MATHUSLA Helo:2018qej ; Curtin:2018mvb ; MATHUSLA:2020uve ; DeVries:2020jbs , could be sensitive to mixing angles $U^{2}$ close to or even reaching the type-I seesaw predictions $U^{2}\sim m_{\nu}/m_{N}$, where $m_{\nu}$ denotes the sum of active neutrino masses and $m_{N}$ the sterile neutrino mass, for $m_{N}$ in the GeV range.
@@
This work investigates the phenomenology of the mLRSM with GeV-scale sterile neutrinos. We perform an explicit matching to the $\nu$SMEFT framework. We then follow the method of Refs. DeVries:2020jbs ; Gunther:2023vmz to investigate the discovery potentials for the light long-lived sterile neutrinos with displaced-vertex (DV) searches to be performed at the LHC and the near detector of the DUNE experiment DUNE:2020lwj ; DUNE:2020ypp ; DUNE:2020jqi ; DUNE:2021cuw ; DUNE:2021mtg of the Long-Baseline Neutrino Facility (LBNF) at Fermilab. For the LHC, [^1] We omit the LHC main experiments ATLAS and CMS, as their usual trigger requirement of a high-$p_{T}$ lepton or jet cannot be satisfied for the sterile neutrinos produced from meson decays as considered here. An exception, though, can arise at the LHCb experiment where specific $B$-parking datasets Bainbridge:2020pgi ; CMS:2024syx could be utilized CMS:2024ita , allowing for probing the sterile-neutrino scenarios studied in this work. Nevertheless, since the $B$-parking searches do not use the full datasets available, we choose not to include this possibility in the work. we focus on the present and future experiments ANUBIS Bauer:2019vqk , FASER and FASER2 Feng:2017uoz ; FASER:2018eoc , MATHUSLA Chou:2016lxi ; Curtin:2018mvb ; MATHUSLA:2020uve , FACET Cerci:2021nlb , CODEX-b Gligorov:2017nwh , and MoEDAL-MAPP1 and MoEDAL-MAPP2 Pinfold:2019nqj ; Pinfold:2019zwp , and the recently approved beam-dump experiment SHiP SHiP:2015vad ; Alekhin:2015byh ; SHiP:2018xqw ; SHiP:2021nfo . Since light sterile neutrinos with RH interactions can significantly impact $0\nu\beta\beta$ rates Bamert:1994qh ; Blennow:2010th ; Mitra:2011qr ; deGouvea:2011zz ; Li:2011ss ; Barry:2013xxa ; Ghosh:2013nya ; Girardi:2013zra ; Barea:2015zfa ; Bolton:2019pcu ; Asaka:2020wfo ; Jha:2021oxl ; Dekens:2024hlz , we also make a detailed comparison to complementary probes provided by rare-meson-decay searches and $0\nu\beta\beta$ experiments. Somewhat remarkably, we find that these very different probes have a comparable reach.
This paper is structured as follows. In Sec. 2, we introduce the mLRSM. In Sec. 3, we detail the $\nu$SMEFT framework. We discuss sterile neutrino production and decay in the mLRSM in Secs. 4 and 5, respectively. We elaborate on the phenomenology of neutrinoless double beta decay in the mLRSM in Sec. 6. In Sec. 7, We explain the numerical simulation and computation procedures for collider analysis based on the theoretical scenarios discussed in the previous sections. We present the numerical results in Sec. 8 and provide our conclusions and an outlook for future work in Sec. 9.
@@ 2 Minimal left-right symmetric model
where all parameters $\kappa,\,\kappa^{\prime},\,v_{L},\,v_{R},\,\alpha,$ and $\theta_{L}$ are real-valued Deshpande:1990ip . Imposing the relation $\sqrt{\kappa^{2}+\kappa^{\prime 2}}=v$, with $v=246$ GeV being the electroweak scale, ensures the correct masses for the SM gauge bosons $Z$ and $W^{\pm}\equiv W_{L}^{\pm}$. Consistency with the absence of right-handed currents in experiments requires the relation $v_{R}\gg v\gg v_{L}$. The phases $\alpha$ and $\theta_{L}$ induce CP violation, and, as elaborated upon below, the vevs $v_{L}$ and $v_{R}$ induce the Majorana neutrino masses.
@@
where since $v_{R}\gg v\gg v_{L}$ one expects $M_{R}\gg M_{D}\gg M_{L}$.
@@
Note that, in comparison with the notations used in Refs. Senjanovic:2011zz ; Nemevsek:2012iq ; Senjanovic:2016vxw , we denote $U_{\text{PMNS}}=V_{L}$ and $U_{R}^{*}=V_{R}$. In the above, the flavor-space projectors $P$ and $P_{s}$ are $6\times 3$ matrices and defined as
@@ 2.1.1 Canonical type-II scenario
In the canonical type-II seesaw scenario, Dirac masses of the neutrinos vanish, which results in the block-diagonal mass matrix
@@
between the active $(m_{1,2,3})$ and sterile neutrino mass eigenstates $(M_{4,5,6})$, irrespective of the choice of the additional discrete LR symmetry. The kinematic accessibility of the two lightest sterile neutrinos, $\nu_{4,5}$ with masses $M_{4,5}$, depends on the choice of hierarchy and the choice of $m_{1}$ and $M_{6}$. The masses of the sterile neutrinos can be expressed in terms of the active-neutrino masses once we fix the highest sterile-neutrino mass and decide the choice of neutrino-mass hierarchy. The relations in Eq. (32) are depicted in Fig. 1.
Neutrino oscillation experiments have measured active-neutrino mixing angles and squared mass differences between the active-neutrino mass eigenstates. We parameterize the PMNS matrix as Workman:2022ynf
@@
The decay rates of sterile neutrinos into SM particles – and vice versa – will depend on the sterile-neutrino masses, RH gauge boson masses, and mixing angles in the diagonalization matrix $U$; see Eq. (11). In the generalized parity symmetry $\mathcal{P}$ scenario, a lower limit on the RH gauge boson masses can be determined through low-energy experimental bounds, independent of the sterile neutrino masses. In principle, measurements of CP-violation in kaon decays strongly limit $M_{W_{R}}\geq 17$ TeV Bertolini:2019out ; Dekens:2021bro . The bound weakens to $M_{W_{R}}\geq 5.5$ TeV when applying an additional Peccei-Quinn mechanism since additional freedom appears in such a non-trivial model extension. This weaker bound is comparable to limits from direct searches at the LHC Dekens:2021bro .
@@ 2.1.2 Canonical type-I Scenario
It is important to emphasize that, unlike the case in the type-II seesaw phenomenology, there are no direct relations between the active- and sterile-neutrino masses. Hence, the sterile-neutrino masses $M_{4,5,6}$ are free parameters in the type-I seesaw scenario and could thus vary over a broad mass range. In principle, $M_{4,5,6}$ could all be kinematically allowed in searches for long-lived HNLs. We provide an overview of independent parameters given the seesaw scenarios under consideration in Tab. 2.
@@ 3.1 The effective neutrino Lagrangian
Since we assume that $v_{R}\gg v_{L}$, additional (charged) Higgs scalars are too heavy to significantly contribute to sterile-neutrino production and decay rates. The relevant terms in the charged-current dim-6 $\nu$SMEFT Lagrangian containing at least one sterile neutrino are
@@
$C^{(6)}_{du\nu l,jk}={V^{R}_{jk}}\>C_{\nu l}^{(6)},\qquad C^{(6)}_{\varphi\nu l}=-\frac{1}{v_{R}^{2}}\frac{2\xi}{1+\xi^{2}}e^{-i\alpha},$ (42)
@@
$\begin{split}\mathcal{L}_{\nu\text{LEFT}}^{(6){\rm CC}}=&\frac{2G_{F}}{\sqrt{2}}\sum_{i,j,l}\Bigg{\{}\sum_{k}\Big{(}\overline{u}_{L}^{i}\gamma^{\mu}d_{L}^{j}\left[C^{(6)}_{\text{VLL1},ijkl}\overline{e}_{L}^{k}\gamma_{\mu}\nu^{l}+C^{(6)}_{\text{VLR1},ijkl}\overline{e}_{R}^{k}\gamma_{\mu}\nu^{l}\right]\\ &+C^{(6)}_{\text{VRR},ijkl}\overline{u}^{i}_{R}\gamma^{\mu}d_{R}^{j}\overline{e}_{R}^{k}\gamma_{\mu}\nu^{l}\Big{)}+\overline{\nu}_{L}^{i}\gamma^{\mu}e^{i}_{L}\left[C^{(6)}_{\text{VLL2},jk}\overline{e}_{L}^{j}\gamma_{\mu}\nu^{l}+C^{(6)}_{\text{VLR2},ijk}\overline{e}_{R}^{j}\gamma_{\mu}\nu^{l}\right]\Bigg{\}},\end{split}$ (43)
where we have rotated to the neutrino mass basis, cf. Eqs. (15) and (16). In the above, $G_{F}=(\sqrt{2}v^{2})^{-1}$ is the Fermi constant, ${u}_{L,R}$ and ${d}_{L,R}$ are chiral up- and down-type quark fields, $e_{L,R}$ are charged lepton fields, and $i,j,k=\{1,2,3\}$ are flavor indices, whereas $l=\{1,\dots,6\}$ denotes the neutrino mass eigenstates. The Wilson coefficients are given by
@@
${C_{\text{VLR1},ijkl}^{(6)}}$ $=V_{ij}v^{2}(P_{s}U^{*})_{kl}\>C^{(6)}_{\varphi\nu l},$ ${C_{\text{VLR2},ijkl}^{(6)}}=\frac{C_{\text{VLR1},ijkl}^{(6)}}{V_{ij}},$ (45)
@@
where $i=\{1\dots 6\}$ is flavor indices for the neutrino mass eigenstates, $j,k=\{1,2,3\}$ are flavor indices for fermion fields $f\in\{u_{L},d_{L},u_{R},d_{R},l_{L},\nu_{L},l_{R}\}$, and
@@ 4.1 Minimal mixing scenario
It is simplest to assume that only the lightest sterile neutrino is kinematically available to be produced from the meson decays. From this point forward, for notational clarity, we will denote the sterile neutrino mass eigenstates as $N_{4},N_{5}$ and $N_{6}$. The sterile-neutrino mass range we consider is $0<M_{4}<m_{B_{s}}\sim 5.37\text{ GeV}$. Unless explicitly stated otherwise, we follow Ref. deVries:2022nyh for our decay-rate calculations, including the phase-space integration, the relevant form factor parameterizations, and the decay constants.
@@ 4.2 Left-right symmetric models
where the meson $M\in\{B,D,K\}$ and the charged lepton $\ell\in\{e,\mu,\tau\}$. Since only the absolute squares of the mixing angles are relevant, there is no dependence on the complex phases $\lambda_{1},\lambda_{2}$, and $\theta_{L}$ that appear in $U_{R}$. The choice of the discrete LR symmetry is also irrelevant to the meson decay rates in this scenario. This greatly simplifies the available parameter space. The choice of the heaviest neutrino mass, $M_{6}$, fixes the mass relation between the active neutrino mass $m_{1}$ and the lightest sterile neutrino mass $M_{4}$ and does not impact the meson decay rates as long as $M_{6}$ is large enough to make $N_{5}$ kinematically unavailable, cf. Eq. (32). For simplicity, we set the CP-violating phase $\alpha=0$. From Tab. 2, in the limit where only $N_{4}$ is kinematically available, it is then apparent that only two relevant benchmark parameters remain. These are the $\xi=\kappa^{\prime}/\kappa$, which determines the $W_{L}$-$W_{R}$ mixing strength, and $M_{W_{R}}$, which is the mass of the RH charged gauge boson.
@@
where $h_{\rm PS}$ and $h_{V}$ are pseudoscalar mesons and vector mesons, respectively, and the minus (plus) sign on the right-hand side of Eq. (51) corresponds to the projection operator $P_{L}$ ($P_{R}$). This relative minus sign in Eq. (51) – which also appears in the two-body leptonic decays – leads to destructive interference terms linear in $\xi$. These effects are sizeable for large $\xi$. Explicitly, the decay rates will be proportional to a factor $|C_{\rm VRR}^{(6)}\mp C_{\rm VLR1}^{(6)}|^{2}$. In the limit where the CP-violating phase $\alpha=0$, we have $C_{\rm VLR1}^{(6)}=\frac{2\xi}{1+\xi^{2}}C_{\rm VRR}^{(6)}$, cf. Eqs. (42), (45) and (46). Hence, for values of $\xi$ approaching 1, the three-body decays into a pseudoscalar meson, and the two-body leptonic decays are suppressed.
@@
For instance, given the lightest sterile neutrino at the pion threshold ($M_{4}=0.14$ GeV), the RH-current contributions will dominate for $M_{W_{R}}\leq 50$ TeV if $m_{1}=1$ meV. Employing the above relation, we see that for relatively light $M_{W_{R}}<10$ TeV and $\xi=0$, type-I seesaw scenarios converge to type-II scenarios since the minimal-mixing contributions are negligible for every reasonable choice of the lightest active-neutrino mass [^2] This requires $\sum\limits_{i=1,2,3}m_{i}<0.12$ eV, considering cosmological constraints Giusarma:2016phn . in the sterile-neutrino mass range $m_{\pi}<M_{4}<m_{B_{s}}$.

### v2

@@ 1 Introduction
Recently, there has been much interest in sterile neutrinos with GeV-scale masses (see, for instance, Refs. Atre:2009rg ; Bondarenko:2018ptm ; Helo:2013esa ; Drewes:2019fou ; Cottin:2022nwp ; Liu:2023gpt ), especially in the context of searches for long-lived particles (LLPs) (see Refs. Alimena:2019zri ; Lee:2018pag ; Curtin:2018mvb ; Beacham:2019nyx for recent reviews of LLPs). In the minimal scenario where the only interactions between the sterile neutrinos and the SM particles are mediated by the active-sterile neutrino mixings $U^{2}$, the current bounds on the sterile neutrinos in this mass range hint that these sterile neutrinos, if existent, would be long-lived Deppisch:2015qwa ; Bryman:2019bjg ; Abdullahi:2022jlv . Beam-dump, neutrino, and collider experiments produce mesons in very large numbers. These mesons can produce sterile neutrinos with masses in the MeV-GeV range through rare decays. With a long lifetime, these sterile neutrinos travel a macroscopic distance before potentially decaying inside a detector volume placed a few to a few hundred meters away. Past studies have investigated these minimal scenarios extensively and indicated that (proposed) future experiments such as CODEX-b Helo:2018qej ; Aielli:2019ivi , DarkQuest Batell:2020vqn , DUNE Krasnov:2019kdc ; Ballett:2019bgd ; Berryman:2019dme ; Coloma:2020lgy ; Moghaddam:2022tac ; Plows:2022gxc ; Gunther:2023vmz , ANUBIS Hirsch:2020klk ; DeVries:2020jbs , SHiP SHiP:2015vad ; Alekhin:2015byh ; SHiP:2018xqw ; DeVries:2020jbs , and MATHUSLA Helo:2018qej ; Curtin:2018mvb ; MATHUSLA:2020uve ; DeVries:2020jbs , could be sensitive to mixing angles $U^{2}$ close to or even reaching the type-I seesaw predictions $U^{2}\sim m_{\nu}/m_{N}$, where $m_{\nu}$ denotes the sum of active-neutrino masses and $m_{N}$ the sterile neutrino mass, for $m_{N}$ in the GeV range.
@@
This work investigates the phenomenology of the mLRSM with GeV-scale sterile neutrinos. [^1] Refs. Li:2022cuq ; Urquia-Calderon:2023dkf studied the phenomenology of long-lived sterile neutrinos with masses of order $10$-$100$ GeV in the mLRSM at, respectively, the LHC and lepton colliders. We perform an explicit matching to the $\nu$SMEFT framework. We then follow the method of Refs. DeVries:2020jbs ; Gunther:2023vmz to investigate the discovery potentials for the light long-lived sterile neutrinos with displaced-vertex (DV) searches to be performed at the LHC and the near detector of the DUNE experiment DUNE:2020lwj ; DUNE:2020ypp ; DUNE:2020jqi ; DUNE:2021cuw ; DUNE:2021mtg of the Long-Baseline Neutrino Facility (LBNF) at Fermilab. For the LHC, [^2] We omit the LHC main experiments ATLAS and CMS, as their usual trigger requirement of a high-$p_{T}$ lepton or jet cannot be satisfied for the sterile neutrinos produced from meson decays as considered here. An exception, though, can arise at the CMS experiment where specific $B$-parking datasets Bainbridge:2020pgi ; CMS:2024syx ; CMS:2024zhe could be utilized CMS:2024ita , allowing for probing the sterile-neutrino scenarios studied in this work. Nevertheless, since the $B$-parking searches do not use the full datasets available, we choose not to include this possibility in the work. we focus on the present and future experiments ANUBIS Bauer:2019vqk , FASER and FASER2 Feng:2017uoz ; FASER:2018eoc , MATHUSLA Chou:2016lxi ; Curtin:2018mvb ; MATHUSLA:2020uve , FACET Cerci:2021nlb , CODEX-b Gligorov:2017nwh , and MoEDAL-MAPP1 and MoEDAL-MAPP2 Pinfold:2019nqj ; Pinfold:2019zwp , and the recently approved beam-dump experiment SHiP SHiP:2015vad ; Alekhin:2015byh ; SHiP:2018xqw ; SHiP:2021nfo . Since light sterile neutrinos with RH interactions can significantly impact $0\nu\beta\beta$ rates Bamert:1994qh ; Blennow:2010th ; Mitra:2011qr ; deGouvea:2011zz ; Li:2011ss ; Barry:2013xxa ; Ghosh:2013nya ; Girardi:2013zra ; Barea:2015zfa ; Bolton:2019pcu ; Asaka:2020wfo ; Jha:2021oxl ; Dekens:2024hlz , we also make a detailed comparison to complementary probes provided by rare-meson-decay searches and $0\nu\beta\beta$ experiments. Somewhat remarkably, we find that these very different probes have a comparable reach.
This paper is structured as follows. In Sec. 2, we introduce the mLRSM. In Sec. 3, we detail the $\nu$SMEFT framework. We discuss sterile-neutrino production and decay in the mLRSM in Secs. 4 and 5, respectively. We elaborate on the phenomenology of neutrinoless double beta decay in the mLRSM in Sec. 6. In Sec. 7, we explain the numerical simulation and computation procedures for collider analysis based on the theoretical scenarios discussed in the previous sections. We present the numerical results in Sec. 8 and provide our conclusions and an outlook for future work in Sec. 9.
@@ 2 Minimal left-right symmetric model
where all parameters $\kappa,\,\kappa^{\prime},\,v_{L},\,v_{R},\,\alpha,$ and $\theta_{L}$ are real-valued Deshpande:1990ip . Imposing the relation $\sqrt{\kappa^{2}+\kappa^{\prime 2}}=v$, with $v=246$ GeV being the electroweak scale, ensures the correct masses for the SM gauge bosons $Z$ and $W^{\pm}\equiv W_{L}^{\pm}$. Consistency with electroweak precision observables Melfo:2011nx and the absence of right-handed currents in experiments require $v_{R}\gg v\gg v_{L}$. The phases $\alpha$ and $\theta_{L}$ induce CP violation, and, as elaborated upon below, the vevs $v_{L}$ and $v_{R}$ induce the Majorana neutrino masses.
@@
where, since $v_{R}\gg v\gg v_{L}$, one expects $M_{R}\gg M_{D}\gg M_{L}$, although this also depends on the choice of Yukawa couplings.
@@
Note that, in comparison with the notations used in Refs. Senjanovic:2011zz ; Nemevsek:2012iq ; Senjanovic:2016vxw , we denote $U_{\text{PMNS}}=V_{L}$ and $U_{R}^{*}=V_{R}$. In the equations above, the flavor-space projectors $P$ and $P_{s}$ are $6\times 3$ matrices and defined as
@@ 2.1.1 Canonical type-II scenario
Dirac Yukawa couplings are assumed to vanish in the canonical type-II seesaw dominance limit, hence $M_{D}=0$. This results in a block-diagonal neutrino mass matrix
@@
between the active- $(m_{1,2,3})$ and sterile-neutrino mass eigenstates $(M_{4,5,6})$, irrespective of the choice of the additional discrete LR symmetry. The kinematic accessibility of the two lightest sterile neutrinos, $\nu_{4,5}$ with masses $M_{4,5}$, depends on the choice of hierarchy and the choice of $m_{1}$ and $M_{6}$. The masses of the sterile neutrinos can be expressed in terms of the active-neutrino masses once we fix the highest sterile-neutrino mass and decide the choice of neutrino-mass hierarchy. The relations in Eq. (32) are depicted in Fig. 1.
Neutrino-oscillation experiments have measured active-neutrino mixing angles and squared mass differences between the active-neutrino mass eigenstates. We parameterize the PMNS matrix as Workman:2022ynf
@@
The decay rates of sterile neutrinos into SM particles – and vice versa – will depend on the sterile-neutrino masses, RH gauge-boson masses, and mixing angles in the diagonalization matrix $U$; see Eq. (11). In the generalized parity symmetry $\mathcal{P}$ scenario, a lower limit on the RH gauge-boson masses can be determined through low-energy experimental bounds, independent of the sterile-neutrino masses. In principle, measurements of CP-violation in kaon decays strongly limit $M_{W_{R}}\geq 17$ TeV Bertolini:2019out ; Dekens:2021bro . The bound weakens to $M_{W_{R}}\geq 5.5$ TeV when applying an additional Peccei-Quinn mechanism since additional freedom appears in such a non-trivial model extension. This weaker bound is comparable to limits from direct searches at the LHC Dekens:2021bro .
@@ 2.1.2 Canonical type-I Scenario
It is important to emphasize that, unlike in the type-II seesaw phenomenology, there are no direct relations between the active- and sterile-neutrino masses. Hence, the sterile-neutrino masses $M_{4,5,6}$ are free parameters in the type-I seesaw scenario and could thus vary over a broad mass range. In principle, $M_{4,5,6}$ could all be kinematically allowed in searches for long-lived HNLs. We provide an overview of independent parameters given the seesaw scenarios under consideration in Tab. 2.
@@ 3.1 The effective neutrino Lagrangian
Since we assume that $v_{R}\gg v_{L}$, additional (charged) Higgs scalars are too heavy to contribute to sterile-neutrino production and decay rates significantly. The relevant terms in the charged-current dim-6 $\nu$SMEFT Lagrangian containing at least one sterile neutrino are
@@
$C^{(6)}_{du\nu l,jk}=-\frac{V^{R}_{jk}}{v_{R}^{2}},\qquad C^{(6)}_{\varphi\nu l}=-\frac{1}{v_{R}^{2}}\frac{2\xi}{1+\xi^{2}}e^{-i\alpha},$ (42)
@@
$\begin{split}\mathcal{L}_{\nu\text{LEFT}}^{(6){\rm CC}}=&\frac{2G_{F}}{\sqrt{2}}\sum_{i,j,l}\Bigg{\{}\sum_{k}\Big{(}\overline{u}_{L}^{i}\gamma^{\mu}d_{L}^{j}\left[C^{(6)}_{\text{VLL1},ijkl}\overline{e}_{L}^{k}\gamma_{\mu}\nu^{l}+C^{(6)}_{\text{VLR1},ijkl}\overline{e}_{R}^{k}\gamma_{\mu}\nu^{l}\right]\\ &+C^{(6)}_{\text{VRR},ijkl}\overline{u}^{i}_{R}\gamma^{\mu}d_{R}^{j}\overline{e}_{R}^{k}\gamma_{\mu}\nu^{l}\Big{)}+\overline{\nu}_{L}^{i}\gamma^{\mu}e^{i}_{L}\left[C^{(6)}_{\text{VLL2},jl}\overline{e}_{L}^{j}\gamma_{\mu}\nu^{l}+C^{(6)}_{\text{VLR2},jl}\overline{e}_{R}^{j}\gamma_{\mu}\nu^{l}\right]\Bigg{\}},\end{split}$ (43)
where we have rotated to the neutrino mass basis, cf. Eqs. (15) and (16). In the equation above, $G_{F}=(\sqrt{2}v^{2})^{-1}$ is the Fermi constant, ${u}_{L,R}$ and ${d}_{L,R}$ are chiral up- and down-type quark fields, $e_{L,R}$ are charged lepton fields, and $i,j,k=\{1,2,3\}$ are flavor indices, whereas $l=\{1,\dots,6\}$ denotes the neutrino mass eigenstates. The Wilson coefficients are given by
@@
${C_{\text{VLR1},ijkl}^{(6)}}$ $=V_{ij}v^{2}(P_{s}U^{*})_{kl}\>C^{(6)}_{\varphi\nu l},$ ${C_{\text{VLR2},kl}^{(6)}}=\frac{C_{\text{VLR1},ijkl}^{(6)}}{V_{ij}},$ (45)
@@
where $i=\{1\dots 6\}$ are flavor indices for the neutrino mass eigenstates, $j,k=\{1,2,3\}$ are flavor indices for fermion fields $f\in\{u_{L},d_{L},u_{R},d_{R},l_{L},\nu_{L},l_{R}\}$, and
@@ 4.1 Minimal mixing scenario
It is simplest to assume that only the lightest sterile neutrino is kinematically available to be produced from the meson decays. From this point forward, for notational clarity, we will denote the sterile neutrino mass eigenstates as $N_{4},N_{5}$ and $N_{6}$. The sterile-neutrino mass range we consider is $0<M_{4}<m_{B_{s}}\sim 5.37\text{ GeV}$. Unless explicitly stated otherwise, we follow Ref. deVries:2022nyh for our decay-rate calculations, including the phase-space integration, the relevant form-factor parameterizations, and the decay constants.
@@ 4.2 Left-right symmetric models
where the meson $M\in\{B,D,K\}$ and the charged lepton $\ell\in\{e,\mu,\tau\}$. Since only the absolute squares of the mixing angles are relevant, there is no dependence on the complex phases $\lambda_{1},\lambda_{2}$, and $\theta_{L}$ that appear in $U_{R}$. The choice of the discrete LR symmetry is also irrelevant to the meson decay rates in this scenario. This greatly simplifies the available parameter space. The choice of the heaviest neutrino mass, $M_{6}$, fixes the mass relation between the active-neutrino mass $m_{1}$ and the lightest sterile-neutrino mass $M_{4}$ and does not impact the meson decay rates as long as $M_{6}$ is large enough to make $N_{5}$ kinematically unavailable, cf. Eq. (32). For simplicity, we set the CP-violating phase $\alpha=0$. From Tab. 2, in the limit where only $N_{4}$ is kinematically available, it is then apparent that only two relevant benchmark parameters remain. These are the $\xi=\kappa^{\prime}/\kappa$, which determines the $W_{L}$-$W_{R}$ mixing strength, and $M_{W_{R}}$, which is the mass of the RH charged gauge boson.
@@
where $h_{\rm PS}$ and $h_{V}$ are pseudoscalar mesons and vector mesons, respectively, and the minus (plus) sign on the right-hand side of Eq. (51) corresponds to the projection operator $P_{L}$ ($P_{R}$). This relative minus sign in Eq. (51) – which also appears in the two-body leptonic decays – leads to destructive interference terms linear in $\xi$. These effects are sizable for large $\xi$. Explicitly, the decay rates will be proportional to a factor $|C_{\rm VRR}^{(6)}\mp C_{\rm VLR1}^{(6)}|^{2}$. In the limit where the CP-violating phase $\alpha=0$, we have $C_{\rm VLR1}^{(6)}=\frac{2\xi}{1+\xi^{2}}C_{\rm VRR}^{(6)}$, cf. Eqs. (42), (45) and (46). Hence, for values of $\xi$ approaching 1, the three-body decays into a pseudoscalar meson, and the two-body leptonic decays are suppressed.
@@
For instance, given the lightest sterile neutrino at the pion threshold ($M_{4}=0.14$ GeV), the RH-current contributions will dominate for $M_{W_{R}}\leq 50$ TeV if $m_{1}=1$ meV. Employing the relation above, we see that for relatively light $M_{W_{R}}<10$ TeV and $\xi=0$, type-I seesaw scenarios converge to type-II scenarios since the minimal-mixing contributions are negligible for every reasonable choice of the lightest active-neutrino mass [^3] This requires $\sum\limits_{i=1,2,3}m_{i}<0.12$ eV, considering cosmological constraints Giusarma:2016phn ; Vagnozzi:2017ovm ; Vagnozzi:2018jhn ; Tanseri:2022zfe ; Planck:2018vyg . in the sterile-neutrino mass range $m_{\pi}<M_{4}<m_{B_{s}}$.